If a videotape cannot be recorded on for the first 1 1/3 feet, and there is a 180 foot video tape, how much can be recorded on by solving the equation: 180= t + 1 1/3 for t.
Answer by stanbon(75887)

180= t + 1 1/3
180 = t+(4/3)
t=180-(4/3)
t=178 2/3
Cheers,
Stan H.